A party rental company has chairs and tables for rent. The total cost to rent 3 chairs and 5 tables is $57. The total cost to rent 12 chairs and 2 tables is $39.
What is the cost to rent each chair and each table?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)


x = price to rent one chair
y = price to rent one table
prices are in dollars

The total cost to rent 3 chairs and 5 tables is $37
It leads to the equation of
3x+5y = 37

Another equation we can form is
12x+2y = 39
because 12 chairs (12x dollars) and 2 tables (2y dollars) cost 39 dollars combined

Let's solve for y in the second equation
12x+2y = 39
2y = 39-12x
y = (39-12x)/2
y = (39/2)-(12x/2)
y = 19.5 - 6x

Then we can plug it into the first equation to solve for x
3x+5y = 57
3x+5(19.5 - 6x) = 57
3x+97.5 - 30x = 57
-27x+97.5 = 57
-27x = 57-97.5
-27x = -40.5
x = (-40.5)/(-27)
x = 1.50
That decimal value is exact.

Then we can use this x value to find y
y = 19.5 - 6x
y = 19.5 - 6(1.50)
y = 19.5 - 9
y = 10.50
That decimal is exact as well

-----------------------------------------

Answers:
One chair costs $1.50 to rent
One table costs $10.50 to rent

Answer by ikleyn(52787)

.
A party rental company has chairs and tables for rent.
The total cost to rent 3 chairs and 5 tables is $57. The total cost to rent 12 chairs and 2 tables is $39.
What is the cost to rent each chair and each table?
~~~~~~~~~~~~~~~~~~~~~~~

Let x be the cost to rent a chair;  

let y be the cost to rent a table.


Write two equations for two unknowns as you read the problem

    3x  + 5y = 57     (1)

    12x + 2y = 39     (2)


This system is good to solve it by Elimination.  For it, multiply equation (1) by 4 (both sides).
Keep equation (2) as is.  Then you have this modified system

    12x + 20y = 4*57      (3)

    12x +  2y =   39      (4)


Now subtract equation (4) from equation (3).  The terms with "x" will delete each other,
and you will get an equation for single unknown "y"

         20y - 2y = 4*57 - 39

Simplify and find y

            18y  = 189

              y  = 189/18 = 10.50  dollars.


Then from equation (2) find 

    12x + 2*10.50 = 39

    12x           = 39 - 21 = 18

     x                      = 18/12 = 1.50 dollars.


ANSWER.  It cost $10.50 to rent one table and $1.50 to rent one chair.


CHECK.   To check, substitute the found values into original equations:

         Eq(1):  3*1.50 + 5*10.50 = 57 dollars,  correct.

         Eq(2): 12*1.50 + 2*10.50 = 39 dollars,  correct.


Solved.

-----------------

To see many other similar and different solved word problems,  look into the lessons
    - Typical word problems on systems of 2 equations in 2 unknowns
    - HOW TO algebreze and solve this problem on 2 equations in 2 unknowns
in this site.

Learn the subject from there.